A boat went up the stream for 1/4 of the distance at 4 mph, then another 1/4 at 2 mph and the rest at 8 mph. What was her average speed?
Answer by RAY100(1637)

velocity = distance / time,,,,,t= d/v
.
1/4 /4 + 1/4 /2 + 1/2 /8 = 1/v avg
.
.0625 + .125 + .0625 = 1/v
.
.25 =1/v
.
v=4 mph
.
checking,,,let Distance = 16 miles (arbitrary)
4/4 + 4/2 + 8/8 = 4 = 16/4 ,,,,,ok